### v1

# AmodalGen3D: Generative Amodal 3D Object Reconstruction from Sparse Unposed Views

Abstract
Reconstructing 3D objects from a few unposed and partially occluded views is a common yet challenging problem in real-world scenarios, where many object surfaces are never directly observed. Traditional multi-view or inpainting-based approaches struggle under such conditions, often yielding incomplete or geometrically inconsistent reconstructions. We introduce AmodalGen3D, a generative framework for amodal 3D object reconstruction that infers complete, occlusion-free geometry and appearance from arbitrary sparse inputs. The model integrates 2D amodal completion priors with multi-view stereo geometry conditioning, supported by a View-Wise Cross Attention mechanism for sparse-view feature fusion and a Stereo-Conditioned Cross Attention module for unobserved structure inference. By jointly modeling visible and hidden regions, AmodalGen3D faithfully reconstructs 3D objects that are consistent with sparse-view constraints while plausibly hallucinating unseen parts. Experiments on both synthetic and real-world datasets demonstrate that AmodalGen3D achieves superior fidelity and completeness under occlusion-heavy sparse-view settings, addressing a pressing need for object-level 3D scene reconstruction in robotics, AR/VR, and embodied AI applications.
1 Introduction
Reconstructing 3D objects from a few unposed and partially occluded images is a fundamental but challenging task in computer vision. In realistic scenarios, objects are often captured from arbitrary viewpoints, with large portions hidden by surrounding structures or other objects. Such incomplete observations make the reconstruction problem inherently amodal: a system must infer the geometry and appearance of regions that are never directly observed. This ability is crucial for object-level 3D scene understanding and downstream applications such as robotic interaction, AR/VR content creation, and embodied AI perception.
Figure 1: Comparison of 3D object reconstruction from sparse, unposed, and occluded inputs. From left to right: input sparse views, results from (a) MVS-based reconstruction (e.g., VGGT [wang2025vggt]), (b) 2D amodal completion lifted to 3D (TRELLIS [xiang2025structured]), and (c) AmodalGen3D (ours). Our generative model faithfully reconstructs complete, occlusion-free 3D geometry consistent with sparse-view constraints.
Traditional multi-view stereo (MVS) and neural rendering approaches, such as NeRF or Gaussian Splatting, rely on dense and accurately calibrated views to estimate intact and consistent geometry. When only a few unposed observations are available, these methods produce incomplete or misaligned reconstructions, as shown in Fig. 1(a). On the other hand, 2D amodal completion techniques can hallucinate missing regions in image space, but their outputs are inconsistent across views, leading to unstable or implausible 3D geometry when combined, as illustrated in Fig. 1(b). The key challenge lies in jointly reasoning about both visible and occluded regions while maintaining geometric and appearance consistency under sparse, unposed constraints.
To address this challenge, we introduce AmodalGen3D, a generative framework for amodal 3D object reconstruction from sparse, unposed views. Our approach treats reconstruction as a conditional generation process that integrates 2D amodal priors with explicit 3D geometric reasoning. Specifically, we first perform object-centric amodal completion on each view to recover both visible and occluded regions. The completed views are then encoded by a DINO-based feature extractor and aggregated through a View-Wise Cross Attention module that maintains spatial consistency across sparsely sampled viewpoints. To infer missing geometry hidden from occluders, we condition the generation on partial geometry estimated from an MVS model, represented as a point cloud. This information is incorporated through a Stereo-Conditioned Cross Attention module equipped with a geometry-guided gating mechanism, which modulates feature fusion based on 3D spatial cues. After processing the generated sparse structure through another pretrained Flow Transformer, the resulting structured latent representation is then decoded by specific decoders to produce a complete, occlusion-free 3D model that respects the sparse-view constraints while plausibly reconstructing the hidden regions. As shown in Fig. 1(c), our method produces coherent and complete 3D reconstructions even when large parts of the object are never observed.
Extensive experiments on both synthetic and real-world datasets demonstrate that AmodalGen3D achieves superior reconstruction quality compared with existing sparse-view and amodal baselines. Our method improves geometric accuracy while maintaining realistic appearance, achieving consistent performance from single-view to multi-view inputs. These results highlight the potential of generative modeling for practical 3D object-level scene reconstruction under real-world occlusion and view sparsity.
The contribution of this work is threefold:
  • We introduce AmodalGen3D, a generative framework for amodal 3D object reconstruction that can recover complete geometry and appearance from sparse, unposed, and partially occluded inputs.
  • We propose the View-Wise Cross Attention that benefits feature aggregation from sparse views, and the novel Stereo-Conditioned Cross Attention for integrating partial MVS geometry into the generation process.
  • Comprehensive experiments validate that AmodalGen3D achieves superior reconstruction fidelity, completeness, and consistency compared with existing sparse-view and amodal reconstruction methods.
2 Related Work
Multi-View Stereo (MVS). Multi-view stereo (MVS) [FurukawaHernandez2015_MultiViewStereoTutorial, Wang2024_LearningBasedMVS_Survey, Yi2020PyramidMVSNet, Yan2020DenseHybridRMVSNet, wang2024dust3r, leroy2024grounding, zhang2024monst3r, yang2025fast3r, zhang2025flare, wang2025pi, keetha2025mapanything] has been studied for decades as a core technique for 3D reconstruction from multiple input views. Recent advances [Ren_2023_HPMMVS, Wu_2024_GoMVS, Santo_2024_MVCPS_NeuS] in dense correspondence learning have enhanced MVS accuracy and robustness by jointly reasoning about geometry and radiance. These methods achieve fine-grained alignment but remain limited when it comes to calibrated settings and fixed depth ranges. To improve scalability and avoid per-scene optimization, feed-forward extensions [Izquierdo_2025_MVSAnywhere, Cabon_2025_MUSt3R, Liu_2024_MVSGaussian, Dong_2024_DepthRangeFreeMVS] introduce transformer-based architectures and specific techniques for zero-shot multi-view reconstruction. These designs collectively enhance scalability and inference efficiency. A complementary line of work [Rojas_2025_HAMSt3R, Peng_2023_GenS, Rich_2024_DIVLoss] explores unposed and unstructured generalization in MVS.
Despite these advances, outputs from MVS methods are often incomplete or noisy (e.g., point clouds), which lack in semantic and appearance consistency for object-level understanding. In contrast, our approach leverages MVS as the geometric prior explicitly: by conditioning the generation model on partial stereo point clouds, AmodalGen3D infers complete, occlusion-free object structures from sparse and unposed inputs, bridging the gap between recent feed-forward MVS models and object-level reconstruction.
Figure 2: Overview of AmodalGen3D. Given sparse images $S$, visibility masks $M_{vis}$, and occlusion masks $M_{occ}$ indicating the occluded object $O$, AmodalGen3D first generates a sparse structure (stage 1) by aggregating multi-view information in Sec. 3.1 and infers the complete geometric structure from the partial stereo point cloud $\mathbf{P}_{O}$ in Sec. 3.2. Once the sparse structure is obtained, we employ a pretrained amodal SLAT Transformer, controlling texture generation with visibility masks $M_{vis}$ and occlusion masks $M_{occ}$, and then generate the structured latent. This structured latent is decoded into an occlusion-free 3D object with high-quality geometry and appearance.
3D Generative Models. Recent advances in 3D generative models have opened new possibilities for structure-aware reconstruction. Methods leveraging Gaussian Splatting [kerbl20233d] and NeRF [mildenhall2021nerf] as the representation excel at novel view synthesis [ye2024no, truong2023sparf], but struggle to segregate object-level entities, hindering object-level 3D scene understanding. Diffusion-based generative models further enhance visual fidelity, controllability, and scalability by leveraging pretrained 2D diffusion priors [huang2024dreamcontrol, ding2024bidirectional, chen2024vp3d] or 3D-aware latent representations [xiang2025structured, long2024wonder3d, zhao2025hunyuan3d, lin2025kiss3dgen, siddiqui2024assetgen]. These methods typically rely on massive optimization iterations or decently posed input images for proper results. Some works attempt to mitigate missing regions using inpainting-augmented generation [chen2024comboverse, han2024reparo, zhou2024layout]. Other scene-level generative pipelines [chen2024single, liu2022towards, huang2025midi, meng2025scenegen, yao2025cast] synthesize entire 3D scenes from a single image, but they rely on carefully curated data and pose-calibrated views, which limits their effectiveness under real-world, unposed scenarios.
In contrast to prior approaches, AmodalGen3D is a generative 3D reconstruction model tailored for sparse, unposed inputs. By fusing generative priors with partial geometric conditioning, it produces reconstructions that are both geometrically faithful and semantically coherent, remaining consistent with sparse-view constraints while plausibly inferring unseen object regions.
Amodal Completion in 2D/3D. Amodal completion aims to infer the hidden or occluded parts of objects beyond visible evidence, which is fundamental for robust scene understanding and reconstruction. Early studies address this through explicit amodal segmentation and completion tasks [LiM_2016_ECCV, Qi_2019_CVPR, Ling_2020_NeurIPS, Mohan_2022_CVPR, Li_2023_ICCV_MUVA], focusing primarily on 2D perception and annotation-based benchmarks. These methods rely on supervised mask prediction or latent shape priors to hallucinate occluded object regions, but lack in strong generative fidelity and geometric consistency.
Generative 2D inpainting models [Rombach_2022_CVPR, yu2023inpaint] have since become common solutions in image-level amodal completion, synthesizing realistic missing content based on surrounding context. Nowadays, researchers further enhance this capacity by introducing specialized amodal reasoning modules [ao2025open, lee2025tuningfreeamodalsegmentationocclusionfree, fan2025multi] or by fine-tuning diffusion models for occlusion-aware reconstruction [ozguroglu2024pix2gestalt, xu2024amodal, lu2025taco]. Extending beyond single-view imagery, multi-view inpainting models [barda2025instant3dit, weber2023nerfiller] attempt to maintain cross-view coherence but often assume calibrated cameras or dense view coverage. More geometry-aware approaches [Li_2022_ECCV, salimi2025geometry, feng2025objfiller, shi2025imfine, Zhan_2024_CVPR] leverage 3D representation priors for scene-level completion; however, they remain limited in handling object-level amodal reconstruction due to alignment errors and weak semantic reasoning. The most relevant work is Amodal3R [wu2025amodal3r], which trains a native 3D generation model from occluded objects, but it is constrained by data scale and limited occlusion reasoning ability.
In sparse and unposed scenarios, 2D inpainting models can recover locally plausible structures yet introduce cross-view inconsistencies that manifest as geometric artifacts when reconstructed to 3D. Our method integrates these powerful 2D amodal priors into a 3D generative framework, enforcing global geometric coherence through the view-wise and stereo-conditioned attention mechanisms. Compared with single-image 3D amodal methods, AmodalGen3D achieves robust reconstruction under severe occlusion, benefiting from multi-view information aggregation and strong cross-view geometric reasoning, while maintaining the realism of image-level completion.
3 Method
For an arbitrary object $O$ in the given $K$ sparse views $S=\{s^{1},s^{2},...,s^{K}\}$ (typically 1-3 views), we can obtain the visibility masks of $O$ and the corresponding text prompt automatically or manually using vision foundation models, e.g., SAM [kirillov2023segment] and Florence 2 [xiao2024florence]. After that, we follow the occlusion analysis and inpainting strategies from [ao2025open] with minor modifications (using the above obtained visibility mask and prompt) to first perform 2D amodal completion across all views. Altogether, we can obtain visual clues of the desired object $O$ at all $K$ input views: the visibility masks $M_{vis}=\{m^{1}_{v},m^{2}_{v},...,m^{K}_{v}\}$ indicating the visible part of $O$; the inconsistent inpainted views of $O$: $I=\{i^{1},i^{2},...,i^{K}\}$; and the occlusion masks $M_{occ}=\{m^{1}_{o},m^{2}_{o},...,m^{K}_{o}\}$ representing the potential occlusion regions.
We illustrate the overall model structure in Fig. 2 using two views as an example. Since geometric completion plays an important role in amodal 3D reconstruction, our design prioritizes the Sparse Structure Transformer, i.e., stage 1. Given object-centric sparse views $I$ processed using 2D amodal completion methods, we apply the proposed View-Wise Cross Attention (Sec. 3.1) to aggregate multi-view features and fuse their latents. The fused representation is further refined through the Stereo-Conditioned Cross Attention (Sec. 3.2) for unobserved structure inference, guided by observed partial stereo geometry and a geometry-aware gating mechanism. The resulting Sparse Structure is decoded and then passed to the pretrained Structure Latent Transformer from [wu2025amodal3r] for Structured Latent generation. Implementation details of the training/inference proxy and the data engine used to construct the object-centric synthetic dataset for model training are described in Sec. 3.3.
Figure 3: The TRELLIS model shows a strong bias and can result in artifact accumulation when conditioned with multiple views.
3.1 View-Wise Cross Attention
We denote the intermediate latent as $z\in\mathbb{R}^{L\times D}$, where $L$ is the number of tokens and $D$ is the hidden dimension of the tokens. In the standard Cross Attention (denoted as $\mathcal{CA}$), the latent $z$ interacts with the encoded image feature $c$,
$c=\texttt{DINO}(i),\quad z^{\prime}=\mathcal{CA}(z,c).$ (1)
Here, DINOv2 [oquab2023dinov2] serves as the image encoder, extracting features for input image $i$. When multiple images are available, existing approaches typically condition on one view at a time, alternating across sampling steps during the generation process. As shown in Fig. 3, this iterative conditioning disrupts consistency across views, often biasing the generation direction towards certain perspectives or causing geometric artifacts accumulation from different views, especially when sparse views differ significantly due to the 2D amodal completion.
To address the inconsistency issue, we introduce the View-Wise Cross Attention, which enables effective aggregation of 2D completion features from sparse views. Unlike the naïve sequential conditioning that alternates between views, our design allows all views to contribute simultaneously at each sampling step, promoting consistent multi-view reasoning and spatial coherence.
For each inpainted view $i^{n}$, we extract its DINO feature $c^{n}$, forming a set of encoded features $C=[c^{1},c^{2},...,c^{K}]$. We then apply Cross Attention in parallel for each view:
$z_{n}^{\prime}=\mathcal{CA}(z,c^{n}),\quad\forall n\in{1,2,\dots,K}.$ (2)
The resulting view-wise latents $Z^{\prime}=[z_{1}^{\prime},z_{2}^{\prime},...,z_{K}^{\prime}]$ are fused into a single latent $z^{\prime}$ through a visibility-weighted average:
$\!\!\!\!z^{\prime}=\frac{1}{K}\sum_{n=1}^{K}w_{n}z_{n}^{\prime},\;w_{n}=\frac{\tau_{n}}{\sum_{j=1}^{K}\tau_{j}},\;\tau_{n}=\frac{m_{v}^{n}}{m_{o}^{n}+m_{v}^{n}}.$ (3)
$\tau_{n}$ represents the visibility ratio, defined as the proportion of possibly visible region in view $i^{n}$. A final LayerNorm is applied after fusion for numerical stability.
In summary, this parallel design preserves view balance, mitigates bias toward specific perspectives, and yields more coherent latent representations across unposed sparse views during the generation process.
Figure 4: A detailed illustration of our proposed View-Wise Cross Attention and Stereo-Conditioned Cross Attention.
3.2 Stereo-Conditioned Cross Attention Module
While the View-Wise Cross Attention enhances multi-view feature fusion, it lacks explicit spatial correspondence between views. To address this, we incorporate geometric cues from a strong multi-view stereo model [wang2025pi], which provides the observed partial 3D geometry as spatial guidance in our Stereo-Conditioned Cross Attention module.
Given sparse observations $S=\{s^{1},s^{2},...,s^{K}\}$, we obtain the stereo point cloud $\mathbf{P}=\{p^{1},p^{2},...,p^{K}\}=\texttt{MVS}(S)$, where each $p^{n}$ corresponds to a point cloud reconstructed from view $s^{n}$ with pixel-level projection. We extract the partial point cloud of the target object by selecting points projected within the visibility masks:
$\mathbf{P}_{O}=\bigcup_{n=1}^{K}p_{v}^{n},\quad p_{v}^{n}=\{\,p\in p^{n}\mid m_{v}^{n}(p)=1\}.$ (4)
The resulting $\mathbf{P}_{O}$ is voxelized into active voxels $\mathbf{V}_{O}$ with grid size $\varphi$, then encoded by the Sparse Structure VAE, and finally transformed into geometric feature tokens using the same patchify and linear layer for transforming input noise:
$\mathbf{V}_{O}$ $=\texttt{Voxelize}(\mathbf{P}_{O},\varphi),\quad f_{\mathbf{V}_{O}}=\mathbf{Enc}(\mathbf{V}_{O}),$ (5)
$c_{\text{geo}}$ $=\texttt{Linear}(\texttt{Patchify}(f_{\mathbf{V}_{O}})).$
We further introduce a Geometry-Guided Gating MLP that modulates attention using the encoded geometry:
$g$ $=\sigma(\text{MLP}_{gating}(c_{\text{geo}})),$ (6)
$\alpha$ $=\texttt{softmax}\!\left(\frac{QK^{\top}}{\sqrt{D}}+\log(g+\varepsilon)\right),$
where $\alpha$ denotes the attention weights computed from query $Q=W_{q}z^{\prime}$ and key $K=W_{k}c_{\text{geo}}$, and $\varepsilon$ is a small constant for numerical stability in logarithmic computation. This design injects explicit geometric cues into the attention computation, improving structural consistency across unposed views and provide spatial guidance for the unobserved part.
Single-View Alternative. When only one input image is available, the partial geometry guidance can instead be derived from monocular depth estimation models [wang2025moge], providing approximate spatial context for single-view amodal reconstruction.
3.3 Data Engine & Training/Inference Proxy
Data Engine. We leverage object-centric 3D datasets from [xiang2025structured] and convert them into watertight meshes using Manifold [huang2018robust]. To simulate realistic occlusions, we generate 3D-consistent occlusion masks by randomly selecting a seed face on each mesh and iteratively expanding to its neighboring faces. This diffusion process continues outward until reaching a randomly sampled coverage ratio, producing spatially coherent and natural occlusion patterns. We then render object views and their corresponding occlusion masks, discarding camera poses with extreme angles or insufficient occlusion. Finally, completed object views are synthesized by treating the occlusion masks as inpainting regions and applying controlled mask distortions for greater diversity (see supplementary for details).
Training Proxy. During training, the number of sparse views per object is randomly sampled from $\{1,2,3,4\}$ for each batch. After sampling, stereo point clouds are generated online based on the selected views. The resulting partial point clouds are centralized and normalized to ensure training stability. Thanks to the synthesized occluded object inputs and visibility masks, modern multi-view stereo models can still recover plausible partial geometry, providing reliable geometric supervision throughout training.
We optimize the model using the standard conditional flow matching (CFM) [lipman2023flowmatchinggenerativemodeling] objective:
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t,z_{0},\epsilon}\|v_{\theta}(z,t)-(\epsilon-z_{0})\|_{2}^{2},$ (7)
where $\theta$ denotes the network parameters, $z_{0}$ and $\epsilon$ are data and noise samples, $t$ is the timestep, and $z=(1-t)x_{0}+t\epsilon$ represents the interpolated state.
Inference Proxy. During inference, our method can process both scene-level and object-level images. For scene-level images, we first generate the stereo point clouds for the entire scene with $S$, and then extract the partial point clouds for each object using the corresponding visibility masks $M_{vis}$. This approach reduces computational overhead when generating multiple objects within a single scene. All other procedures remain consistent with the training process. Fig. 2 shows an example of scene-level inference.
4 Experiments
Figure 5: Amodal 3D object reconstruction results on the GSO dataset. We give examples generated with 1, 2, and 4 views, respectively. Note that for FreeSplatter [xu2025freesplatter], we adopt the Hunyuan-MV  [yang2024tencent] to generate multiple views under the input of single image (first two rows).
4.1 Experimental Settings
Dataset. We use 3D-FUTURE [fu20213d] and Amazon Berkeley Objects (ABO) [collins2022abo] for 3D consistent occluded object simulation. 3D-FUTURE contains 9472 objects, and ABO contains 4485. We train our network with all objects combined. For each object, we render 150 views while filtering out camera elevation angles lower than -75 degrees or higher than 75 degrees, which results in around 90 views per object that are close to real-world scenarios and captures. We select views whose occluded regions take up from 20% to 60% during training. For evaluation, we use the Google Scanned Object (GSO) [downs2022google] dataset for object-level test. In addition, we sample from Hypersim [roberts2021hypersim] dataset to simulate in-the-scene scenraios; COCO [lin2015microsoftcococommonobjects] and Mip-NeRF 360 [barron2022mipnerf360] dataset to evaluate our in-the-wild performance in Sec. 4.4.
Implementation Details. We make the entire sparse structure generation stage (stage 1) of the model trainable and use 8 RTX 6000 GPUs in training. We use $\pi^{3}$ [wang2025pi] as the multi-view stereo model and normalize the coordinates of partial point clouds. Then we perform voxelization with a grid size of $64^{3}$. We adopt the AdamW [loshchilov2017decoupled] optimizer with a learning rate of 5e-5, and Classifier-Free-Guidance [ho2022classifier] with a drop rate of 0.1 during training. We train the model for 12 epochs with a global batch size of 32, which takes approximately 16 hours to complete. Please refer to the supplementary for more details in the training, inference, and data engine settings.
Figure 6: Additional qualitative results, together with occluded input images after 2D amodal completion.
4.2 Main Results
Comparison Methods We compare our method with various 3D object generation methods equipped with the latest 2D completion methods. First, to validate the 3D amodal completion ability with unposed sparse views of AmodalGen3D, we employ TRELLIS [xiang2025structured], Amodal3R [wu2025amodal3r], and FreeSplatter [xu2025freesplatter] for comparison. We utilize [ao2025open] to perform 2D amodal completion at sparse views. Second, to demonstrate that AmodalGen3D is also effective and can generate high-quality 3D assets in real-world scenarios, we show in-the-wild and in-the-scene qualitative results, compared with Amodal3R [wu2025amodal3r] and SceneGen [meng2025scenegen], respectively.
Figure 7: Visual comparisons from the ablation studies showing the impact of each proposed model design.
Figure 8: In-the-wild 3D amodal reconstruction results. Single-view examples are from the COCO [lin2015microsoftcococommonobjects] dataset; sparse-view examples are from the Mip-NeRF 360 [barron2022mipnerf360] (cropped for better visual layout).
Quantitative Results. We present quantitative results tested on the GSO dataset. We first render 3D consistent masks and then generate all 1030 objects one by one. For a fair comparison, we select those views whose visible parts occupy from 30% to 70% of the whole. To validate the effectiveness and generalizability of our AmodalGen3D, we compare our method with others on different numbers of sparse view settings, i.e., 1, 2, and 4 views. As shown in Tab. LABEL:tab:_exp_metric, we employ Fréchet Inception Distance (FID) [heusel2017gans], Kernel Inception Distance (KID) [binkowski2018demystifying] to evaluate the quality of rendered images; CLIP-Score [radford2021learning] to measure the feature similarity; Minimum Matching Distance (MMD) and Coverage (COV) [nichol2022point] to evaluate the geometry quality.
[…]
Ours full 32.12 5.50 39.08
- Gating MLP 32.87+0.55 5.58+0.08 37.68 -1.40
- View-Wise CA 38.64+6.42 5.76+0.26 38.34 -0.74
- Stereo-Cond. CA 42.92+10.80 5.91+0.41 35.26 -3.82
- all proposed 46.23+14.11 5.94+0.44 34.42 -4.66
Table 1: Ablation study on different proposed modules. Red indicators represent performance drops compared to the full model.
Qualitative Results. We here present qualitative results on the GSO dataset with simulated occlusions for amodal 3D object reconstruction. As shown in Fig. 5, our model outperforms other 3D generation methods with better geometry consistency and coherence. Moreover, with the strong 2D prior and our proposed design, AmodalGen3D shows better potential in uncommon and complex amodal object completion. We showcase more visualization results in Fig. 6, together with the 2D amodal completion results as a better demonstration of AmodalGen3D’s ability in fixing image-level structural inconsistencies in the 3D space.
View Num. FID $\downarrow$ CLIP-Score(%) $\uparrow$ MMD(‰) $\downarrow$
1 33.91 82.23 5.52
2 32.12 -1.79 82.33 +0.10 5.50 -0.02
4 30.73 -1.39 82.53 +0.20 5.48 -0.02
10 29.56 -1.17 82.98 +0.45 5.45 -0.03
20 29.43 -0.13 82.95 -0.03 5.45 0.00
Table 2: Ablation study on AmodalGen3D’s ability with number of input views, we test up to 20 views. The red, green indicator represents performance gain or loss compared to previous one.
2D Amodal Comp. FID $\downarrow$ CLIP-Score(%) $\uparrow$ COV(%) $\downarrow$
pix2gestalt [ozguroglu2024pix2gestalt] 38.61 82.07 37.26
Flux inpainting [flux2024] 33.07 84.26 39.49
OAAC [ao2025open] 32.12 82.33 39.08
Table 3: Ablation study on different 2D Amodal Completion methods. We present results generated with 2 views.
4.3 Ablation Study
Structure Design. We conduct ablation studies to assess the effectiveness of different module designs. Tab. 1 reports quantitative results, and Fig. 7 shows visual comparisons. With View-Wise Cross Attention, our method performs notably better under sparse views, demonstrating better rendering quality with more reasonable sparse structure as prior for texture generation. Moreover, incorporating Stereo-Conditioned Cross Attention for explicit geometry guidance yields a great benefit in the geometric metrics (MMD). As shown in Fig. 7, MVS conditioning helps AmodalGen3D produce more plausible and structurally consistent geometry, even when the 2D completions are inconsistent or contain artifacts at image level.
Number of Views. We investigate the impact of the number of input views on our method. As shown in Tab. 2, the FID, CLIP-Score, and MMD metrics remain stable or exhibit slight improvements, indicating that our method can effectively handle dense settings with up to 20 input images, which demonstrates the adaptability of our work.
2D Amodal Completion. We conduct experiments on how our model performs with different 2D amodal completion methods. As shown in Tab. 3, we can see that for different methods, our AmodalGen3D is able to produce consistent and robust results that are comparable in numerical metrics. Additionally, it is intuitive that stronger completion methods produce better results, which shows the potential and adaptability of our AmodalGen3D with more advanced techniques in the future.
Errors in MVS. In recent MVS models, the reconstructed point clouds occasionally exhibit misaligned or failing structures. Therefore, we analyze how such MVS errors propagate and affect our generation pipeline. As shown in Fig. 9, when the reconstructed stereo is geometrically implausible, our model can be less influenced by the stereo conditioning and perform self-correction with the 3D generative prior.
Figure 9: We demonstrate that AmodalGen3D can resist a certain degree of errors in the MVS reconstructed point clouds, e.g., misalignment and overlapping geometry from different views.
Figure 10: In-the-scene 3D amodal reconstruction results. Input images are from the Hypersim [roberts2021hypersim] dataset. Red and Green icons indicate different target objects (marked in the first row).
4.4 Reconstruction In-the-wild & In-the-Scene
We further evaluate our model in both in-the-wild and in-the-scene settings to assess its robustness and generalization ability. The in-the-wild cases correspond to real-world environments with complex interactions between objects, while the in-the-scene setting mainly involves structured indoor layouts, such as furniture arrangements or tabletop scenes. As shown in Fig. 8, our AmodalGen3D achieves realistic 3D amodal reconstruction across various in-the-wild conditions. In Fig. 10, we show in-the-scene reconstruction results compared with  [wu2025amodal3r] and  [meng2025scenegen], indicating the potential of our model in object-level scene reconstruction under the sparse-view and severe occlusion setting. These results demonstrate that our method can generalize across different domains and retain visible partial information more effectively while faithfully hallucinating unseen parts.
5 Conclusion
In this paper, we propose AmodalGen3D, a 3D object generator capable of reconstructing amodal 3D objects from arbitrary sparse and unposed observations. Using only object-level occlusion training data and simulated 2D amodal completion, AmodalGen3D can seamlessly integrate with 2D amodal completion methods and multi-view stereo models, enabling generalization across both synthetic and real-world scenarios and paving the way for high-quality 3D generation in diverse applications. While our amodal reconstruction relies on priors learned from generative models to infer missing regions, which may look plausible but not always match the ground-truth geometry, our approach still represents a meaningful step forward in amodal 3D reconstruction, demonstrating the potential of combining generative priors with sparse, real-world–level observations.
[…]
Supplementary Material
A More Implementation Details
We give more implementation details on the Network Architecture (Sec. A.1) and Data Engine (Sec. A.2) in this section. Moreover, we present a comprehensive and detailed evaluation setting for our quantitative and qualitative results.
A.1 Network Architecture
In this section, we specify some designs of the proposed network structure.
In the Stereo-Conditioned Cross Attention, we train this module from scratch and adopt the Xavier initialization [glorot2010understanding]. We normalize the point cloud coordinates to [-0.5, 0.5], and then use a grid size (voxel size) of $1/64$ for point cloud voxelization. Moreover, we perform certain noisy point cloud filtering during this process: for each point $p$ in the point cloud, we apply a neighborhood search within a fixed spatial radius of $1/64$, and if there are fewer than 5 points within this radius range, we simply discard the outlier point to achieve cleaner partial point clouds.
The Geometry-Guided Gating MLP is consisted of: 1) the first linear layer that compresses the latent channels to $D/2$, followed by a ReLU activation function; 2) another linear layer that unifies the attention score, followed by a Sigmoid activation function.
Since the TRELLIS model uses a default 12 sampling steps for generation, for multi-view image conditions that are not divisors of 12, we increase the number of sampling steps to meet their minimum requirements.
A.2 Data Engine
We provide additional details for the 3D consistent mask simulation part mentioned in Sec. 3.3. When diffusing 3D consistent masks, we give the range of coverage ratio as $20\%-60\%$. We denote the final generated 3D consistent mask as $M_{3d}$. To synthesize the occlusion masks used for training, we first render the occlusion and visibility parts of each object with the simulated $M_{3d}$, forming $m_{3docc}$ and $m_{3dvis}$. After that, we generate the bounding rectangle (box) $b_{occ}$ for the $m_{3docc}$ and take the pixels inside as the initial canvas. The final occlusion mask is generated by performing erosion 131-3 times with a kernel size of $k=13$, and simultaneously setting $m_{3dvis}$ as all zeros to keep the visibility of simulated objects.
A.3 Evaluation setting
Image-level Quality & Similarity We evaluate the image-level quality of the renderings using we employ Fréchet Inception Distance (FID) [heusel2017gans], Kernel Inception Distance (KID) [binkowski2018demystifying] as mentioned in the main paper. Additionally, we use CLIP-Score [radford2021learning] with ViT-B/32 model to calculate the feature-level cosine similarity between the rendered images and the input images. Specifically, for each object, we render 8 views using camera poses of a constant radius (2) and Field-of-View (40°), with azimuth angles of {0°, 45°, 90°, 135°, 180°, 225°, 270°, 315°} and a constant elevation angle of 15°. We calculate the score for each pair of corresponding views (generated and ground truth), and finally do a mean average for the metric score calculation.
Geometry Quality Moreover, we use Coverage (COV) and Minimum Matching Distance (MMD) with Chanmfer Distance [nichol2022point] to evaluate the geometry quality. We sample 4096 points from each ground truth and generated point cloud, which are obtained from the depth maps from different views using the farthest point sampling (FPS).
B Additional Results
In this section, we present more results on the 3D amodal reconstruction results at single-object level in Sec. B.1 (additional results on the GSO dataset); we also give more in-the-wild and in-the-scene results in Sec. B.2.
For the efficiency evaluation, our AmodalGen3D can typically generate a single 3D object within 15 seconds, which causes a slight computation overhead compared to TRELLIS [xiang2025structured] and Amodal3R [wu2025amodal3r]. However, our method leads to much better reconstruction results, which is worth the increased inference time.
B.1 Additional Single-Object Results
Fig. 11 and Fig. 12 show more examples on the single-object level reconstruction results. We present results with 1, 2, and 4 views, validating the effectiveness of our method. We still compare with TRELLIS [xiang2025structured] and Amodal3R [xu2024amodal] in this part to show that our method can fix inconsistencies and infer high-quality unobserved structures.
B.2 Additional In-the-wild & In-the-scene Results
We give more in-the-wild and in-the-scene examples in Fig. 13 and Fig. 14, showcasing that AmodalGen3D achieves superior fidelity and completeness under various occlusion conditions in real-world scenarios.
Figure 11: More qualitative results with different sparse-view settings (1, 2, and 4 views). We conduct comparisons with TRELLIS and Amodal3R on the Google Scanned Object dataset, demonstrating the effectiveness of our method.
Figure 12: We present more examples generated by AmodalGen3D on the GSO dataset.
Figure 13: Additional in-the-wild amodal 3D object reconstruction results on the MipNeRF 360 [barron2022mipnerf360] dataset.
Figure 14: Additional in-the-scene amodal 3D object reconstruction results on the Hypersim [roberts2021hypersim] dataset.